Use a simple model: a total of N molecules are contained in a box of side length d. Each molecule has a mass m and a velocity u. By considering the collisions between the molecules and the walls of the box, using the assumptions in kinetic theory of gas derive the equation:

pV = 1⁄3 N m<c^2>

where p is the pressure of the gas, V is its volume and <c^2> is the mean square speed of the gas. 


1
Expert's answer
2021-04-01T18:42:23-0400

If the length of the cube, mass of the molecule and velocity are l, m and v respectively,

Momentum in the x-direction "mu_x"

Momentum in the -x-direction "-mu_x"

Change in momentum "2mu_x"

Total time taken - from one end to the other and vice versa "\frac{2l}{u_x}"

Rate of change in momentum "\frac{mu_x^2}{l}"

According to Newton's Second Law, the rate of change of momentum is the force exerted by the molecule on the wall.

Therefore, Force "\frac{mu_x^2}{l}"

Pressure on the wall, "\frac{mu_x^2}{l^3}"

If there are N molecules in the container, "P=\frac{Nmu_x^2}{V}"

Since velocity in each direction is equal,


"P=\frac{Nmc^2}{3V}"